OpenStudy (anonymous):

Set up the integral to find the area between the two curves y=11-7x-x^2 and y=1-4x.

OpenStudy (anonymous):

hey dolphin

OpenStudy (anonymous):

integrate top curve minus bottom curve dx

OpenStudy (anonymous):

looks like top curve is y = 11 - 7x - x^2 , and bottom is 1-4x, better check that

OpenStudy (anonymous):

Thanks cantorset. That's what I thought. I found the limits are from -5 to 2 Should I integrate from -5 to 2 or split it between -5 to 0 and 0 to +2?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

solve 1-4x = 11 - 7x x^2

OpenStudy (anonymous):

yes, I solved and found x=-5 and x=+2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

343/6

OpenStudy (anonymous):

thats the area

OpenStudy (anonymous):

Thanks